OpenStudy (anonymous):

Solve the system 2x + 3y = 3 and 3x - 2y = 11 by using graph paper or graphing technology. What is the solution to the system?

OpenStudy (lucaz):

if you put the equations in y-intercept form you find the y-int and slope y = mx + b, m is the slope and b the y-int. \[\large{2x+3y=3~->~y=\frac{-2}{3}x+1}\] \[\large{3x-2y=11~->~y=\frac{3}{2}x-\frac{11}{2}}\]

OpenStudy (lucaz):

different slopes, the lines intersect at some point, they are actually the negative reciprocal of each other, wich means they are perpendicular lines. the first slope is -2/3 so when x increases by 3, y decreases by 2 the second slope is 3/2 when x increases by 2, y increases by 3 plot the lines.. the intersection point is the solution of the system

OpenStudy (lucaz):